OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x))=X and g(f(x))=x f(x)=(x+4)/6 and g(x)=6x-4

OpenStudy (xapproachesinfinity):

well you got f(g(x))=f(6x-4)=[(6x-4)+4]/6 you get [6x-4+4]/6=6x/6=x same way with the other one